A disk is originally spinning about its center at 4.8 rad/s clockwise. It makes 3.6 revolutions before it is stopped. If under a constant angular acceleration, how long does it take to stop in seconds?
Answer #1

w0 = 4.8 rad/s

theta = (3.6)(2pi) = 22.62 rad

wf = 0

Applying wf^2 - w0^2 = 2 alpha theta

0^2 - 4.8^2 = 2(alpha)(22.62)

alpha = - 0.5093 rad/s^2


wf = w0 + alpha t

0 = 4.8 - 0.5093 t

t = 9.4 sec .............Ans